OpenStudy (anonymous):

Put the equation in slope-intercept form and identify the slope of the line. 9x + 6y = 3 my answer: -1/3

OpenStudy (anonymous):

compare equation with y=mx+c where m is slope and c is intercept

OpenStudy (anonymous):

its x=1/3

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=(-9x+3)/6 2y=-3x+1 {y=mx+c}

OpenStudy (anonymous):

slope is -3/2